Show that Log9xy = 1/2Log3x + 1/2Log3y


1
Expert's answer
2021-08-30T17:18:32-0400

1/2Log3x+1/2Log3y=1/2(Log3x+Log3y)=1/2(Log3(x*y))=Log32(xy)=Log9xy

that show equation is true